# The essential regularity of singular connections in Geometry

### v3

@@ Abstract.
In this paper we accomplish three things: (i) We discover the geometric (true) regularity of affine connections, their essential (highest possible) regularity, a geometric property independent of starting atlas. (ii) We give a checkable necessary and sufficient condition for determining whether or not connections are at their essential regularity, based on the relative regularity of the connection and its Riemann curvature. (iii) We introduce a computable procedure for lifting any $L^{p}$ affine connection in an atlas ($p>n$), to a new atlas in which the connection exhibits its essential regularity. To accomplish this, we prove that the RT-equations, originally designed by the authors to locally lift the regularity of singular connections by just one derivative, surprisingly, also induce an implicit hidden regularization of the Riemann curvature tensor, together with a global regularization of transition maps between regularizing coordinate charts. From this we deduce a multi-step regularization of the connection, and construct a new atlas in which the connection exhibits its essential regularity. The present paper is a culmination of the theory of the RT-equations which provides a computable iterative procedure for lifting an atlas to a new atlas in which the connection exhibits its essential regularity, applicable to any $L^{p}$ affine connection defined in a $W^{2,p}$ starting atlas, $p>n$. Taken all together, this provides a definitive theory for determining whether or not singularities in an $L^{p}$ affine connection are essential or removable by coordinate transformation, together with an explicit procedure for lifting removable singularities to their essential regularity, both locally and globally, $p>n$. This includes GR shock wave singularities and cusp singularities (continuous metrics with infinite gradients), both obstacles to numerical simulation, but not yet black hole type singularities where $p\leq n$ and metrics are unbounded and discontinuous. It is well known that manifolds by themselves do not carry any intrinsic level of regularity. We propose the essential regularity of a connection to be the point where an intrinsic level of regularity enters the subject of Geometry.
@@ 1. Introduction
The problem of determining whether a singularity is removable by coordinate transformation, and determining the true regularity of removable singularities, has been a major open problem in Differential Geometry. In this paper we resolve this problem for any affine connection in $L^{p}$, $p>n$. The question whether a singularity is essential or removable became a central problem in General Relativity (GR) when Eddington and Finkelstein first discovered an explicit coordinate transformation which regularized the singularity present at the event horizon of the black hole in the Schwarzschild metric. Since that time, every reader is left wondering how one could determine at the start that the singularity is removable. This question extends to other important singularities in GR, including singularities in the Kerr and Reissner-Nordstrom metric, singularities at GR shock waves [17, 37, 14, 26, 24, 4], as well as infinite gradient singularities like those that arise in numerical simulations of binary black hole mergers [1, 2, 23]. In fact, the question how to tell, from a map alone, whether a singularity, like the North Pole in the Mercator map, is a problem with the map or a real singularity in the object which is being mapped, is as old as the subject of Differential Geometry. Historically, determining that a singularity is removable has required the explicit construction of a regularizing coordinate transformation, and failing this, neither a definitive test, nor an explicit procedure for constructing regularizing coordinate transformations, has been available for determining apriori whether the search for a regularization is feasible or futile [9, 15, 17, 18, 26, 40]. This problem extends to all levels of regularity, and more fundamentally raises the following question: What is the true (highest possible) regularity a metric, or more generally a connection, can exhibit in an atlas of coordinate charts, and what defines and characterizes the true, highest possible, geometric level of regularity of a connection?
@@
In this paper we accomplish three things: (i) We make precise the notion of the highest possible regularity of a connection, its essential regularity, and establish its consistency as a geometric (true) property of any affine connection defined in an atlas, cf. Definition 2.2. (ii) We give a necessary and sufficient condition, based on the relative regularity of a connection and its Riemann curvature tensor, for when a connection has reached its essential regularity in a given atlas, cf. Theorem 2.5. (iii) We introduce an iterative procedure, starting from any atlas, for constructing a new atlas in which the essential regularity of a connection is achieved, cf. Theorem 2.3. We establish (i) - (iii) in the Sobolev spaces $W^{s,p}$, $s\geq 0$, $p>n$, for any affine connection whose components are $L^{p}$ regular in some starting atlas on a general $n$-dimensional manifold $\mathcal{M}$. Connection regularity $L^{p}$, $p>n$, includes all Hölder and Lipschitz continuous metrics in the case of metric connections, thus addressing GR shock wave solutions constructed by the Glimm scheme as well as cusp singularities with infinite metric gradients, both obstacles to numerical simulation, but not yet the discontinuous and unbounded metrics associated with black hole singularities. [^1] By Christoffel’s formula a metric is always exactly one derivative more regular than its connection. So a metric connection in $L^{p}$ has a metric in $W^{1,p}$. When $p>n$ such metrics are continuous by Morrey’s inequality, but can have infinite gradient; when $p\leq n$ metric components can be discontinuous and unbounded, like the singularity at the event horizon of the Schwarzschild metric. The ideas in this paper potentially extend to other regularity classes and geometries, including connections on vector bundles [32].
@@
To accomplish results (i) - (iii) in this paper, we prove that the RT-equations, originally designed by the authors to locally lift the regularity of singular connections by one derivative, surprisingly, also induce an implicit hidden regularization of the Riemann curvature tensor, together with a global regularization of transition maps between regularizing coordinate charts. This is a major breakthrough beyond authors’ prior results on local one-step regularizations in [29, 30, 31, 33], and we view this paper as a culmination of our work which began with the discovery of the RT-equations.
@@
The RT-equations, introduced by the authors’ in [27, 28, 29, 30, 31, 32, 33], are a non-invariant coordinate dependent system of partial differential equations (PDE) which is elliptic regardless of metric signature, and whose solutions provide the local coordinate transformations which lift an $L^{p}$ connection by one derivative of regularity when its Riemann curvature is in $L^{p}$. This extends the optimal regularity results of Kazdan-DeTurck from Riemannian to non-Riemannian geometry. See also [3, 8] for results in Lorentzian geometry subject to additional assumptions. Several applications of the RT-equations have been established so far, including the problem of existence and uniqueness of geodesic curves at low regularity [34], (cf. [20, 35, 36]), Penrose’s Strong Cosmic Censorship Conjecture [25], (cf. [22, 12, 13, 19]), Quantum Gravity [39], and the first extension of Uhlenbeck compactness [40] to vector bundles over Lorentzian and non-Riemannian manifolds with both compact and non-compact Lie groups [32], (cf. [7, 33, 41]).
@@
In our prior results, we viewed the curvature regularity as fixed, which was sufficient for a one-step regularization by the RT-equations, but not for the multi-step regularization required in general to obtain the essential regularity of a connection, because multi-step regularizations always require a regularization of the curvature as well. The RT-equations do not provide any such explicit regularization of the curvature, and for this reason authors thought for a long time that multi-step regularizations by the RT-equations are not possible. But, in this paper, by leveraging the notion of essential regularity–a consistent notion of regularity not formulated before–, we prove that an implicit (hidden) regularization of the curvature by the RT-equations takes place in each step until the essential regularity of the connection is reached. Moreover, our prior results were only sufficient to obtain a regularization in a neighborhood of any point, but this was not sufficient to regularize a geometry globally, because even when covering a manifold by regularizing coordinate charts, it seemed impossible to control the regularity of the transition maps. This problem is resolved in the present paper by discovering a correlation between Jacobian regularity and the regularity of connections a Jacobian maps between. [^2] After writing this paper we found that Myers-Steenrod [21], Calabi-Hartman [6] and Taylor [38] used similar ideas to control Jacobian regularity of transformations that preserve regularity of Riemannian metrics. McCann et al. extended this to Lorentzian metrics in [5]. We refine this result to a necessary and sufficient condition and to transformations between affine connections of different regularities (cf. Lemma 2.1), and this refinement is crucial for the proofs in this paper. Using this regularity control of Jacobians, together with a consistent notion of essential regularity and a refined existence theory for the RT-equations, we here prove that the RT-equations regularize any connection, together with its curvature, whenever the connection is below its essential regularity; and we prove that regularized connections in different coordinate charts are always related by transition maps which exhibit sufficient regularity to preserve essential regularity, thereby globalizing our methods to a consistent atlas on the whole manifold. This then establishes the existence of a new connection-dependent atlas in which a connection exhibits its essential regularity, a global regularization of an entire geometry to its essential regularity for any $L^{p}$ connection defined in a starting atlas, $p>n$.
@@
Turning to the problem of regularizing singularities, to clarify the ideas, consider first the simplest case of point singularities. The essential regularity of a connection at a point is the highest possible regularity one can obtain in some coordinate chart containing that point (cf. Definition 5.1); and by a point singularity we mean a connection which is below its essential regularity in every restriction of some given coordinate chart to a neighborhood of the point singularity. In this case, the results in this paper imply that a necessary and sufficient condition for a connection $\Gamma$ given in a coordinate chart to be regularizable around a point, (i.e., for a point singularity to be removable), is that the Riemann curvature tensor ${\rm Riem}(\Gamma)={\rm Curl}(\Gamma)+[\Gamma,\Gamma]$ be at least as regular as the connection in some restriction of the chart to some neighborhood of the point, due to a cancellation of singular terms in the curvature; and given the singularity is removable, solving the RT-equations successively until the curvature is one derivative less regular than the connection, provides an explicit iterative procedure for constructing the coordinate transformations which lift $\Gamma$ to its essential (highest possible) regularity in some new coordinate chart containing the point. [^3] One can make an analogy here with the classical problem of determining whether or not a vector field $F$ is conservative: Given the components of a vector field $F$ in some coordinate system, a necessary and sufficient condition for there to exist a potential in a simply connected domain, is the condition $CurlF\equiv 0$; and given $F$ is appropriately Curl-free, the method of partial integration provides a procedure for explicitly computing its potential. Authors’ prior work addressed local one-step regularizations of point singularities, but singularities are generally global, that is, points or surfaces of regularity defects could be scattered, potentially densely, throughout the manifold, like Glimm scheme based GR shock waves. Based on our new global multi-step regularization of singular connections and transition maps by the RT-equations, and the global notion of essential regularity established in Definition 2.2, we obtain the following complete characterization of local (point) and global singularities.
@@
The hierarchy of apparent singularities: A connection is two or more derivatives below its essential regularity if and only if the curvature is at least one derivative more regular than the connection; a connection is precisely one derivative below its essential regularity if and only if the curvature has the exact same regularity as the connection; a connection is at its essential regularity if and only if the curvature is precisely one derivative less regular than the connection, (cf. Theorem 2.5).
@@
Thus, an apparent (removable) singularity is one where the curvature is at least as regular as the connection, and an essential (non-removable) singularity is one where the curvature is precisely one derivative less regular than the connection. [^4] One typically only refers to something as “singular” if its regularity is too low for some construction. Again, our current theory establishes this principle rigorously for regularity measured in the Sobolev spaces $W^{s,p}$, $s\geq 0$, $p>n$. Based on this hierarchy of singularities, we prove in Theorem 2.3 that any apparent singularity can be regularized to essential connection regularity by iterative use of the RT-equations, both locally and globally. To authors’ knowledge, this is the first definitive theory of apparent singularities applicable to Lorentzian geometry and general affine connections. Moreover, because a manifold by itself does not have enough structure to determine a geometric level of regularity, (every $C^{1}$ atlas is compatible with a $C^{\infty}$ atlas [16]), the essential regularity of a connection is the point at which a geometric level of regularity enters the subject of Geometry.
@@
The outline of the paper is as follows: In Section 2 we state and discuss the main results of this paper, including a discussion of $C^{\infty}$ structures. In Section 3 we establish an improved existence theory for the RT-equations, as required for this paper. In Section 4 we prove regularity control of coordinate transformations when connections are mapped between different levels of regularity. Combining the results of Sections 3 and 4, we develop in Section 5 the local versions of the results stated in Section 2. In Section 6 we complete the theory by extending the local results to global ones, thereby completing the proofs of the theorems stated in Section 2. In Section 7 we establish Sobolev estimates for one- and two-step regularizations of connections on compact manifolds. [^5] It is well-known that Sobolev norms in non-Riemannian geometry are in general non-invariant, that is, their values have no invariant meaning. However, Sobolev spaces themselves have invariant meaning within a suitable regular atlas.
@@ 2. Statement of results
Regularizing the components of a connection by coordinate transformation requires measuring the regularity in some space of differentiability, like $C^{k}$, $H^{s}$, $C^{k,\alpha}$ or $W^{s,p}$. In this paper we address affine connections and measure differentiability in the Sobolev spaces $W^{s,p}$, the space of functions with $s$ derivatives integrable in $L^{p}$ [11]. We are interested in the regularity of singular connections on bounded sets, so with slight abuse of terminology we define a function on an unbounded domain in $\mathbb{R}^{n}$ to be in $L^{p}$ if its $L^{p}$ norm on every bounded subset is finite, (essentially we use $L^{p}$ to mean $L^{p}_{loc}$). For simplicity we restrict the theory to non-negative integer values $s\in\mathbb{N}_{0}$ (including $s=0$) and fixed values of $p$, assuming $n<p<\infty$. [^6] The ideas here are more general, and analogous results should in principal carry over to other spaces amenable to elliptic regularity theory applied to the RT-equations, e.g., $W^{s,p}$ for continuous values of $s\geq 0$. The ideas here extend to connections on vector bundles by the results in [32]. The assumption $p>n$ and $s\in\mathbb{N}_{0}$, (and $p>4$ if $n\leq 3$ and $s=0$), is required by authors’ current existence theory for the RT-equations in [31] on which the results in this paper are based. Note also that the case $p=\infty$ is a singular case of elliptic regularity theory, and when $\Gamma$ and ${\rm Riem}(\Gamma)$ are in $L^{\infty}$ the RT-equations give a regularization to $W^{1,p}$ for any $p<\infty$. An affine connection $\Gamma$ defined on $(\mathcal{M},\mathcal{A})$, a manifold $\mathcal{M}$ with atlas $\mathcal{A}$, is said to have regularity $W^{s,p}$, and we write $\Gamma\in W^{s,p}_{\mathcal{A}}$, if the components of $\Gamma$ have regularity $W^{s,p}$ in every coordinate patch $(x,\Omega)$ of $\mathcal{A}$, and for components of $\Gamma$ in a fixed coordinate chart $(x,\Omega)$ we write $\Gamma_{x}\in W^{s,p}(\Omega_{x})$ with $\Omega_{x}\equiv x(\Omega)\subset\mathbb{R}^{n}$. Similarly, we write ${\rm Riem}(\Gamma)\in W^{s,p}_{\mathcal{A}}$ to express that the components of the Riemann curvature tensor have regularity $W^{s,p}$ in each coordinate system $(x,\Omega)$ of the atlas $\mathcal{A}$, and ${\rm Riem}(\Gamma_{x})\in W^{s,p}(\Omega_{x})$ denotes the regularity of curvature components in a fixed coordinate chart $(x,\Omega)$. We let $\mathcal{A}_{s}$ denote an atlas which preserves $W^{s,p}$ connection regularity. Our first lemma establishes that the regularity of a connection controls the regularity of an atlas as follows:
@@
Our necessary and sufficient condition for essential regularity asserts that $\Gamma\in W^{s,p}_{\mathcal{A}}$ exhibits its essential regularity $s=m\equiv ess_{\mathcal{M}}(\Gamma)\in\mathbb{N}_{0}$ if and only if the regularity of ${\rm Riem}(\Gamma)$ is precisely one derivative below the regularity of $\Gamma$ in the defining atlas $\mathcal{A}$; namely, ${\rm Riem}(\Gamma)\in W^{s-1,p}_{\mathcal{A}}$, but ${\rm Riem}(\Gamma)\notin W^{s,p}_{\mathcal{A}}$. This is the case $s=m$ recorded in the following theorem which can be viewed as a globalization of the existence theory for the RT-equations.
@@
    $\Gamma\in W^{s+1,p}_{\mathcal{A}_{s+1}}$ in some subatlas $\mathcal{A}_{s+1}$ of the maximal $W^{2,p}$ atlas $\mathcal{A}^{\text{max}}_{s}$ of $\mathcal{A}_{s}$ if and only if ${\rm Riem}(\Gamma)\in W^{s,p}_{\mathcal{A}_{s}}$.

### v4

@@ Abstract.
This paper, a culmination of the authors’ theory of the RT-equations, accomplishes the following: (i) We discover there is a true (geometric) regularity associated with every affine connection, its essential regularity, the highest possible regularity achievable by coordinate transformation, a geometric property independent of starting atlas. (ii) We give a checkable necessary and sufficient condition for determining whether or not a connection is at its essential regularity in a given atlas, based on the relative regularity of the connection and its Riemann curvature. (iii) We introduce a computable procedure based on the RT-equations for lifting any $L^{p}$ affine connection given in a starting atlas, to a new atlas in which the connection exhibits its essential regularity. This resolves the long-standing problem of determining whether or not a singularity in an affine connection is removable or essential, applicable to any connection with components locally in $L^{p}$, $p>n$, general enough to include GR shock wave and cusp singularities in General Relativity. Since a manifold by itself does not carry an intrinsic level of regularity, the authors propose that the essential regularity of a connection marks the point at which an intrinsic level of regularity enters the subject of geometry.
@@ 1. Introduction
The problem of determining whether or not a singularity is removable by coordinate transformation, and the true regularity of removable singularities, has been a central open problem in Differential Geometry and General Relativity from the beginning. In this paper we resolve the problem for any affine connection locally in $L^{p}$ in a given atlas, $p>n$. The question whether a singularity is essential or removable became central to General Relativity (GR) when Eddington and Finkelstein first discovered an explicit coordinate transformation which regularized the singularity present at the event horizon of the black hole in the Schwarzschild metric. Since that time, every reader is left wondering how one could determine at the start that the black hole singularity is removable. This question extends to other important singularities in GR, including singularities in the Kerr and Reissner-Nordstrom metrics, singularities at GR shock waves [17, 39, 14, 27, 25, 4], as well as infinite gradient singularities like those that arise in numerical simulations of binary black hole mergers [1, 2, 24]. In fact, the question how to tell, from a map alone, whether a singularity, like the North Pole in the Mercator map, is a problem with the map or a real singularity in the object which is being mapped, is as old as the subject of Differential Geometry. Historically, determining that a singularity is removable has required the explicit construction of a regularizing coordinate transformation, and failing this, neither a definitive test, nor an explicit procedure for constructing regularizing coordinate transformations, has been available for determining a priori whether the search for a regularization is feasible or futile [9, 15, 17, 18, 27, 41]. This problem extends to all levels of regularity, and more fundamentally raises the following question: What is the true, highest possible, regularity a metric, or more generally a connection, can exhibit in an atlas of coordinate charts, and what defines and characterizes this true, highest possible, geometric level of regularity of a connection?
@@
In this paper we accomplish three things: (i) We make precise the notion of the highest possible regularity of a connection, its essential regularity, and establish its consistency as a geometric property of any affine connection, independent of starting atlas, cf. Definition 2.2. (ii) We give a necessary and sufficient condition, based on the relative regularity of a connection and its Riemann curvature tensor, for when a connection has reached its essential regularity in a given atlas, cf. Theorem 2.5. (iii) We introduce an iterative procedure based on the RT-equations, starting from any atlas, for constructing a new atlas in which the essential regularity of a connection is achieved, cf. Theorem 2.3. We establish (i) - (iii) in the Sobolev spaces $W^{s,p}_{\rm loc}$, $s\geq 0$, $p>n$, for any affine connection whose components are locally $L^{p}$ regular in some starting atlas on a general $n$-dimensional manifold $\mathcal{M}$. Connection regularity $L^{p}$, $p>n$, includes all Hölder and Lipschitz continuous metrics in the case of metric connections, thus addressing GR shock wave solutions constructed by the Glimm scheme as well as cusp singularities (continuous metrics with infinite gradients), both obstacles to numerical simulation, but not yet the discontinuous and unbounded metrics associated with black hole singularities. [^1] By Christoffel’s formula a metric is always exactly one derivative more regular than its connection. So a metric connection in $L^{p}$ has a metric in $W^{1,p}$. When $p>n$ such metrics are continuous by Morrey’s inequality, but can have infinite gradient; when $p\leq n$ metric components can be discontinuous and unbounded, like the singularity at the event horizon of the Schwarzschild metric. The ideas in this paper potentially extend to other regularity classes and geometries, including connections on vector bundles [33].
@@
To accomplish results (i) - (iii) in this paper, we prove that the RT-equations, originally designed by the authors to locally lift the regularity of singular connections by one derivative, surprisingly, also induce an implicit hidden regularization of the Riemann curvature tensor, together with a global regularization of transition maps between regularizing coordinate charts. This is a culmination of the authors’ prior results on local one-step regularizations which began with the discovery of the RT-equations.
@@
The RT-equations, introduced by the authors’ in [29] and analysed in [30, 31, 32, 34], are a non-invariant coordinate dependent system of partial differential equations (PDE) which is elliptic regardless of metric signature, and whose solutions provide the local coordinate transformations which lift an $L^{p}$ connection by one derivative of regularity when its Riemann curvature is also in $L^{p}$. This extends the optimal regularity results of Kazdan-DeTurck from Riemannian to non-Riemannian geometries. See also [3, 8] for results in Lorentzian geometry subject to additional assumptions. Several applications of the RT-equations have been established so far, including the problem of existence and uniqueness of geodesic curves at low regularity [35], (cf. [21, 37, 38]), Penrose’s Strong Cosmic Censorship Conjecture [26], (cf. [23]), Hawking’s Singularity Theorem [19], (cf. [12, 13, 20]), and the first extension of Uhlenbeck compactness [41] to vector bundles over Lorentzian and non-Riemannian manifolds allowing for compact as well as non-compact Lie groups [33], (cf. [7, 34, 42]). The authors’ research program began with the problem of singular GR shock waves, constructed in coordinates where Lorentzian metrics are below their essential regularity [4, 14, 17, 27, 28, 39].
@@
In our prior results we viewed the curvature regularity as fixed, which was sufficient for a one-step regularization by the RT-equations, but not for the multi-step regularization required to extend our theory to a consistent definition of the essential regularity of a connection and the coordinate transformations which achieve it, because multi-step regularizations always require a regularization of the curvature as well. The RT-equations do not provide any such explicit regularization of the curvature, and for this reason authors thought for a long time that multi-step regularizations by the RT-equations are not possible. But, in this paper, by leveraging the notion of essential regularity–a consistent notion of regularity not formulated before–we prove that an implicit (hidden) regularization of the curvature by the RT-equations takes place in each step until the essential regularity of the connection is reached. Moreover, our prior results were only sufficient to obtain a regularization in a neighborhood of any point, but this was not sufficient to regularize a geometry globally, because even when covering a manifold by regularizing coordinate charts, it seemed impossible to control the regularity of the transition maps. This problem is resolved in the present paper by discovering a correlation between Jacobian regularity and the regularity of connections a Jacobian maps between. [^2] Myers-Steenrod [22], Calabi-Hartman [6] and Taylor [40] used similar ideas to control Jacobian regularity of transformations that preserve regularity of Riemannian metrics. McCann et al. extended this to Lorentzian metrics in [5]. We refine this result to a necessary and sufficient condition and to transformations between affine connections of different regularities (cf. Lemma 2.1), and this refinement is crucial for the proofs in this paper. Using this regularity control of Jacobians, together with a consistent notion of essential regularity and a refined existence theory for the RT-equations, we prove that the RT-equations regularize both a connection and its curvature by one derivative, iteratively, until the connection reaches its essential regularity. Moreover, we prove that connections regularized up to their essential regularity, starting from different coordinate charts, always generate transition maps which exhibit the requisite regularity required to preserve the essential regularity, thereby extending essential regularity to a consistent definition applicable to any global atlas in which connection components lie locally in $L^{p}$, $p>n$. Taken together, this establishes for any connection given in an atlas the existence of a new atlas in which the connection exhibits its essential regularity, thus achieving a global regularization of an entire geometry to its essential regularity, applicable to any $L^{p}$ connection defined in any starting atlas of regularity $W^{2,p}$ or higher, $p>n$.
@@
Turning to the problem of regularizing singularities, to clarify the ideas, consider first the simplest case of point singularities. The essential regularity of a connection at a point is the highest possible regularity one can obtain in some coordinate chart containing that point (cf. Definition 5.1); and by a point singularity we mean a connection which is below its essential regularity in every restriction of some given coordinate chart to a neighborhood of the point singularity. In this case, the results in this paper imply that a necessary and sufficient condition for a connection $\Gamma$ given in a coordinate chart to be regularizable around a point, (i.e., for a point singularity to be removable), is that the Riemann curvature tensor ${\rm Riem}(\Gamma)={\rm Curl}(\Gamma)+[\Gamma,\Gamma]$ be at least as regular as the connection in some restriction of the chart to some neighborhood of the point, due to a cancellation of singular terms in the curvature; and given the singularity is removable, solving the RT-equations successively until the curvature is one derivative less regular than the connection, provides an explicit iterative procedure for constructing the coordinate transformations which lift $\Gamma$ to its essential (highest possible) regularity in some new coordinate chart containing the point. [^3] One can make an analogy here with the classical problem of determining whether or not a vector field $F$ is conservative: Given the components of a vector field $F$ in some coordinate system, a necessary and sufficient condition for there to exist a potential in a simply connected domain, is the condition $CurlF\equiv 0$; and given $F$ is appropriately Curl-free, the method of partial integration provides a procedure for explicitly computing its potential. The authors’ prior work addressed local one-step regularizations of point singularities, but singularities are generally global, that is, points or surfaces of regularity defects could be scattered, potentially densely, throughout the manifold, like Glimm scheme based GR shock waves. Based on our new global multi-step regularization of singular connections and transition maps by the RT-equations, and the global notion of essential regularity established in Definition 2.2, we obtain the following complete characterization of local (point) and global singularities.
@@
The hierarchy of apparent singularities: A connection is two or more derivatives below its essential regularity if and only if the curvature is at least one derivative more regular than the connection; a connection is precisely one derivative below its essential regularity if and only if the curvature has the exact same regularity as the connection; a connection is at its essential regularity if and only if the curvature is exactly one derivative less regular than the connection, (see Theorem 2.5 for the precise statement). [^4] The proof requires the theory of the RT-equations for non-Riemannian connections. In the case of positive definite Riemannian metric connections, the methods of Kazdan-DeTurck appear to suffice in place of the RT-equations, but present authors are not aware of a statement of this result in the literature for Riemannian metric connections.
@@
Thus, at levels of regularity low enough to be considered problematic, an apparent (removable) singularity is one where the curvature is at least as regular as the connection, and an essential (non-removable) singularity is one where the curvature is precisely one derivative less regular than the connection. Again, our current theory establishes this principle rigorously for regularity measured in the Sobolev spaces $W^{s,p}_{\rm loc}$, $s\geq 0$, $p>n$. Based on this hierarchy of singularities, we prove in Theorem 2.3 that any apparent singularity in a connection can be removed by regularizing the connection to its essential regularity by iterative use of the RT-equations, both locally and globally. To the authors’ knowledge, this is the first definitive theory of apparent singularities applicable to Lorentzian geometry and general affine connections.
Even though Sobolev norms in non-Riemannian geometry are in general not invariant, we establish here that the highest possible level of Sobolev regularity, the essential regularity, is invariant. In [36], the authors argue that the essential regularity of the gravitational metric, determined experimentally by large length-scale measurements, would encode information about the underlying physics on smaller length scales. Because a manifold by itself does not have enough structure to determine a geometric level of regularity, (every $C^{1}$ atlas is compatible with a $C^{\infty}$ atlas [16]), the essential regularity of a connection marks the point at which a geometric level of regularity enters the subject of geometry.
@@
The outline of the paper is as follows: In Section 2 we state and discuss the main results of this paper, including a discussion of $C^{\infty}$ structures. In Section 3 we establish an improved existence theory for the RT-equations, as required for this paper. In Section 4 we prove regularity control of coordinate transformations when connections are mapped between different levels of regularity. Combining the results of Sections 3 and 4, we develop in Section 5 the local versions of the results stated in Section 2. In Section 6 we complete the theory by extending the local results to global ones, thereby completing the proofs of the theorems stated in Section 2. In Section 7 we establish Sobolev estimates for one- and two-step regularizations of connections on compact manifolds.
@@ 2. Statement of results
Regularizing the components of a connection by coordinate transformation requires measuring the regularity in some space of differentiability, like $C^{k}$, $H^{s}$, $C^{k,\alpha}$ or $W^{s,p}$. In this paper we measure differentiability in the Sobolev spaces $W^{s,p}$, the space of functions with $s$ derivatives integrable in $L^{p}$ with standard Lebesgue measure using the Euclidean metric of the coordinate charts in $\mathbb{R}^{n}$. Differentiation is then taken with respect to partial derivatives in $\mathbb{R}^{n}$. For example, given a function $f$ on an open set $\Omega\subset\mathbb{R}^{n}$, we write $f\in W^{s,p}(\Omega)$ if $f\in L^{p}(\Omega)$ and with partial derivatives $\partial_{x}^{k}f\in L^{p}(\Omega)$ for all k=1,,sk=1,...,s, [11]. We are interested in the regularity of singular connections on bounded sets, so with slight abuse of terminology we define a function on an unbounded domain in $\mathbb{R}^{n}$ to be in $L^{p}$ if its $L^{p}$ norm on every bounded subset is finite, (essentially, we use $L^{p}$ to mean $L^{p}_{\rm loc}$ and $W^{s,p}$ for $W^{s,p}_{\rm loc}$). For clarity we restrict the theory to non-negative integer values $s\in\mathbb{N}_{0}$ (including $s=0$) and fixed values of $p$, assuming $n<p<\infty$. [^5] The ideas here are more general, and analogous results should in principal carry over to other spaces amenable to elliptic regularity theory applied to the RT-equations, e.g., $W^{s,p}$ for continuous values of $s\geq 0$. The ideas here extend to connections on vector bundles by the results in [33]. The assumption $p>n$ and $s\in\mathbb{N}_{0}$, (and $p>4$ if $n\leq 3$ and $s=0$), is required by the authors’ current existence theory for the RT-equations in [32] on which the results in this paper are based. Note also that the case $p=\infty$ is a singular case of elliptic regularity theory, and when $\Gamma$ and ${\rm Riem}(\Gamma)$ are in $L^{\infty}$ the RT-equations give a regularization to $W^{1,p}$ for any $p<\infty$. In this paper we address affine connections defined on a manifold $\mathcal{M}$ with atlas $\mathcal{A}$. Given a coordinate chart $(x,\Omega)$ in $\mathcal{A}$, we let $\Gamma_{x}$ denote connection components represented in $x$-coordinates, and we let $\Gamma$ denote the collection of all representations $\Gamma_{x}$ of $\mathcal{A}$, i.e., $\Gamma\equiv\big\{\Gamma_{x}\mid(x,\Omega)\in\mathcal{A}\big\}$. For simplicity always refer to $\Gamma$ as an (affine) connection; note, (though never used in this paper), in our notation $\nabla_{x}\equiv\partial_{x}+\Gamma_{x}$ defines a covariant derivative on the tangent bundle of $\mathcal{M}$ represented in $x$-coordinates. An affine connection $\Gamma$, defined on $(\mathcal{M},\mathcal{A})$, is said to have regularity $W^{s,p}$, if the components of $\Gamma$ have regularity $W^{s,p}$ in every coordinate patch $(x,\Omega)$ of $\mathcal{A}$, in which case we write $\Gamma\in W^{s,p}_{\mathcal{A}}$, and for components of $\Gamma$ in a fixed coordinate chart $(x,\Omega)$ we write $\Gamma_{x}\in W^{s,p}(\Omega_{x})$ with $\Omega_{x}\equiv x(\Omega)\subset\mathbb{R}^{n}$. Similarly, we write ${\rm Riem}(\Gamma)\in W^{s,p}_{\mathcal{A}}$ to express that the components of the Riemann curvature tensor have regularity $W^{s,p}$ in each coordinate system $(x,\Omega)$ of the atlas $\mathcal{A}$, and ${\rm Riem}(\Gamma_{x})\in W^{s,p}(\Omega_{x})$ denotes the regularity of curvature components in a fixed coordinate chart $(x,\Omega)$. We let $\mathcal{A}_{s}$ denote an atlas which preserves $W^{s,p}$ connection regularity, (so $\mathcal{A}_{s}$ can be any atlas with transition maps of regularity $W^{2,p}$ or higher). [^6] It is well-known that Sobolev norms are typically non-invariant in non-Riemannian geometry, i.e., their values vary with the coordinate system or depend on additional structures, like a Riemannian background metric. In the context of this paper such a background metric can only be maintained in $W^{1,p}$, a regularity too low to be of any use. Sobolev spaces, on the other hand, have invariant meaning in the sense that if a tensor or connection is in $W^{s,p}$ in one coordinate chart, then it will also be in $W^{s,p}$ in any other coordinate chart within any atlas of sufficiently regular transition maps, $W^{s+1,p}$ for tensors, and $W^{s+2,p}$ for connections. Our first lemma establishes that the regularity of a connection controls the regularity of an atlas as follows:
@@
Our necessary and sufficient condition for essential regularity asserts that $\Gamma\in W^{s,p}_{\mathcal{A}}$ exhibits its essential regularity $s=m\equiv ess_{\mathcal{M}}(\Gamma)\in\mathbb{N}_{0}$ if and only if the regularity of ${\rm Riem}(\Gamma)$ is precisely one derivative below the regularity of $\Gamma$ in the defining atlas $\mathcal{A}$; namely, ${\rm Riem}(\Gamma)\in W^{s-1,p}_{\mathcal{A}}$, but ${\rm Riem}(\Gamma)\notin W^{s,p}_{\mathcal{A}}$. This is recorded in the following more general theorem, which establishes boundedness of the curvature (in suitable Sobolev spaces) as a definitive necessary and sufficient condition which determines when a (singular) connection can be regularized.
@@
    $\Gamma\in W^{s+1,p}_{\mathcal{A}_{s+1}}$ in some subatlas $\mathcal{A}_{s+1}$ of the maximal $W^{2,p}$ atlas $\mathcal{A}^{\text{max}}_{s}$ of $\mathcal{A}_{s}$ if and only if ${\rm Riem}(\Gamma)\in W^{s,p}_{\mathcal{A}_{s}}$. [^7] Note that $\mathcal{A}_{s+1}$ actually lies within the smaller atlas $\mathcal{A}_{s}^{\text{max}(s+2)}$, because the transition maps between $\mathcal{A}_{s}$ and $\mathcal{A}_{s+1}$ preserve connection regularity $W^{s,p}$, and by Lemma 2.1, have regularity $W^{s+2,p}$. The subatlas $\mathcal{A}_{s+1}$ constructed via the RT-equations always lies in $\mathcal{A}_{s}^{\text{max}(s+2)}\subset\mathcal{A}^{\text{max}}_{s}$, (cf. Section 6).